# Simulations in Statistical Workflows

### v1

@@ 4.2 Parameter Estimation
Posterior distributions: For a given prior $\pi(\theta)$, we are interested in replacing the prior with the posterior $\pi(\theta\mid y)$ which incorporates all information about $\theta$ carried by the data $y$,
@@
Needless to say, both problems are computationally challenging for non-analytic models (e.g., no closed-form for the Bayesian posterior or unknown frequentist sampling distribution). The situations is further aggravated by P models defined solely through a (randomized) algorithm for generating data $y$ from parameter configurations $\theta$ without explicitly assuming a parametric data model $\pi(y\mid\theta)$ (Diggle and Gratton,, 1984). For such implicit models, the construction of confidence sets is challenging not only because we cannot evaluate the likelihood, but also because we need to test null hypotheses across the entire parameter space (Dalmasso et al.,, 2024). By the same token, posterior approximation is difficult because not only the marginal likelihood $\pi(y)$ involves an intractable integral, but also the implicit likelihood (Cranmer et al.,, 2020), rendering posterior approximation doubly intractable.
@@
Despite their theoretical elegance (Frazier et al.,, 2018), most ABC methods share a fundamental limitation: computations must be repeated from scratch each time a model is fit to new data. Yet, there are many scenarios in which multiple model refits are necessary—for example, when modeling multiple data sets (von Krause et al.,, 2022), performing in silico model verification (Section 3), or aiming for real-time inference (Zeng et al.,, 2025). To address these challenges, a different way of utilizing simulations was needed—a way to pool or “compile” computations into global estimators that can produce near-instant results for arbitrary queries (Gershman and Goodman,, 2014; Paige and Wood,, 2016; Le et al.,, 2017). In hindsight, neural networks, as modular and flexibly composable universal function approximators, proved to be the ideal choice.
@@
The core idea of amortized inference (Bürkner et al.,, 2023; Zammit-Mangion et al.,, 2024) is to leverage simulations as training data for neural networks that learn to approximate any quantity of interest, such as point estimates, test statistics, or full posterior distributions (cf. Figure 5). We refer to this process as simulation-based training.
@@
^$\widehat{\mathcal{R}}(y_{\text{obs}}):=\{\theta\in\Theta\mid\lambda(y_{\text{obs}};\theta)\geq\widehat{C}_{\alpha}\}.$ (8)
@@
$\mathcal{L}(q):=\mathbb{E}_{\pi(\theta,y)}\Big{[}\mathcal{S}(q(\theta\mid y),\theta)\Big{]}\approx\frac{1}{M}\sum_{m=1}^{M}\mathcal{S}(q(\theta\mid y^{(m)}),\theta^{(m)}),$ (9)
which would guarantee $q(\theta\mid y)=\pi(\theta\mid y)$ under perfect convergence for large simulation budgets $M\rightarrow\infty$ (Gneiting et al.,, 2007) and universal density approximators $q$ (Draxler et al.,, 2024). For instance, using the log-score for $S$ (Pacchiardi et al.,, 2024), we retrieve the popular maximum likelihood objective,
@@
Equation 10 requires generative models $q(\theta\mid y)$ with tractable densities, such as normalizing flows. More recently, score-based diffusion models have entered ABI (Sharrock et al.,, 2022; Geffner et al.,, 2023; Gloeckler et al.,, 2024), aimed at estimating the score of the (log) posterior density $\nabla_{\theta}\log\pi(\theta\mid y)$ by minimizing a tractable re-formulation of the weighted Fisher divergence:
@@
where $w(t)$ is a positive weight function, $s$ is a neural network, and $\pi_{t}(\theta^{(t)},y)$ is the model distribution defined in diffusion time $t$. The most popular tractable version of Equation 11 is conditional denoising (Song et al.,, 2020), but the framework includes various other “free-form” model families capable of learning from simulations, such as flow matching (Albergo et al.,, 2023; Lipman et al.,, 2022).
Notably, it was not initially evident that neural networks could serve as standalone global posterior approximators. Early approaches, for example, combined neural networks with ABC (Blum and François,, 2010) or SMC (Paige and Wood,, 2016). Papamakarios and Murray, (2016, p. 3) noted that generative networks could be trained over the entire prior predictive distribution $\pi(y)$, but dismissed the approach as “grossly inefficient”. However, we now understand that amortization is not wasteful, but even crucial for efficiently generating model-based insights across fields as disparate as cognitive science (von Krause et al.,, 2022) and astrophysics (Dax et al.,, 2025), among others.
Importantly, the probabilistic calibration of all amortized methods can be efficiently checked via SBC (see Section 3.1 and Section 5.2), since the cost of independent posterior sampling over multiple model simulations is amortized. Additionally, the closeness of observed data to the typical set of simulations can be assessed via out-of-distribution (OOD) detection techniques in (embedded) data space (Schmitt et al., 2023a, ; Huang et al.,, 2023).
Amortized methods underscore a fruitful merger between scientific simulation and deep learning. They also suggest a roadmap for future research addressing some of the most pressing limitations for the field. This includes the search for robust amortized methods that integrate both simulated and real data, while maintaining high simulation efficiency in particularly challenging settings with expensive simulators, limited data, and high dimensionality (Elsemüller et al.,, 2025; Mishra et al.,, 2025).
@@ 5 Model Checking
The final workflow step of model checking investigates the compatibility of the statistical model with the observed data after inference. This data-conditional perspective differs conceptually from model verification (see Section 3): While model verification typically builds on simulated data sets (e.g., from the prior predictive distribution), model checking focuses on the actual observed data set $y_{\text{obs}}$ to gauge the inference validity for the data of interest. However, simulations still play an important role for scaling model checking algorithms to complex, high-dimensional use cases.
@@ 5.1 Predictive Checking
If the data generating process is formalized as a simulation program, we can synthetically generate data replications based on the inferred parameter values. Predictive checking (Gelman and Hill,, 2006) combines parameter estimation (inverse problem) with simulation (forward problem), thereby casting the inference results back into the data space.
@@
In the Bayesian framework, we can propagate all uncertainty from the posterior distribution through the data model. The result is the posterior predictive distribution, which expresses the probability distribution of new data $y^{\prime}$ while accounting for all uncertainty in the probabilistic system,
@@
$\theta^{(s)}\sim\pi(\theta\mid y_{\text{obs}}),\quad y^{\prime(s)}_{1},\ldots,y^{\prime(s)}_{N}\sim\pi(y\mid\theta^{(s)})\quad\text{for}\;s=1,\ldots,S,$ (14)
where each simulated data set $y^{\prime(s)}=\{y^{\prime(s)}_{1},\ldots,y^{\prime(s)}_{N}\}$ contains $N$ observations. The draws from the posterior predictive distribution represent uncertainty-aware replications of the observed data (see Figure 6 for an illustrative example), where all associated uncertainty is propagated through the forward simulation process. Akin to frequentist predictive checking, any discrepancy statistic can then be computed from the observed data and the distribution of replicated data. For example, one might check whether the observed maximum value or a certain autocorrelation is extreme relative to the distribution of those metrics across the simulated replicates. If the observed data exhibit features that would be very unlikely under the posterior predictive distribution, it indicates a model misfit.
@@
Figure 6: Statistical modeling of global temperature data with three parameters. Based on the observed global temperature data $y_{\text{obs}}$ (6(a)), Bayesian inference yields draws from the posterior distribution of the statistical model parameters $\theta$ (6(b)). These posterior draws can be re-inserted into the forward simulation program (i.e., the global temperature model $f:\theta\mapsto y$) to obtain draws from the posterior predictive distribution (6(c)). These data replications are again global temperatures and contain the propagated uncertainty from the initial inference procedure.
@@ 5.2 Posterior SBC
As detailed in Section 3.1.1, simulation-based calibration (SBC) checks if the inference is well-calibrated for data simulated from the prior predictive distribution. However, after observing real data $y_{\text{obs}}$, it is often more relevant to assess the inference conditional on that specific data rather than the entire prior predictive space. To address this, Posterior SBC (Säilynoja et al.,, 2025) uses the SBC principle to validate the model implementation and inference algorithm conditioned on the observed data through means of simulations. This approach first draws $S$ samples $\theta^{\prime(s)}\sim\pi(\theta\mid y_{\text{obs}})$ from the (approximate) posterior conditional on the observed data $y_{\text{obs}}$. For each posterior draw, we then simulate data replications $y^{\prime(s)}$ from the posterior predictive distribution according to Equation 14. Finally, for each simulated data replication $y^{\prime(s)}$, we draw $D$ augmented posterior samples $\theta^{\prime\prime(s,1)},\ldots,\theta^{\prime\prime(s,D)}$ conditional on the observed data $y_{\text{obs}}$ and the respective data replication $y^{\prime(s)}$,
@@
and finally check whether the “old” posterior $\pi(\theta\mid y_{\text{obs}})$ and the augmented posterior $\pi(\theta\mid y_{\text{obs}},y^{\prime})$ are well-calibrated. Pursuing a similar data-driven perspective, Fazio et al., (2024) propose adding small portions of real or simulated data to improve the robustness of Bayesian simulations, which in turn leads to increased stability for SBC and model simulations more broadly.
@@ 5.3 Model comparison
where $\pi(M)$ is a discrete prior distribution over the candidate models. In prior-based model comparison, the central quantity is the marginal likelihood $\pi(y_{\text{obs}}\mid M_{l})=\int\pi(y_{\text{obs}}\mid\theta,M_{l})\,\pi(\theta\mid M_{l})\mathrm{d}\theta$, which quantifies the evidence for model $M_{l}$ by integrating over all parameter values. This enables the computation of Bayes Factors (Kass and Raftery,, 1995) and posterior model probabilities. Since computing the marginal likelihood requires solving a potentially high-dimensional integral, this family of model comparison methods is typically computationally infeasible for complex models. Based on synthetic simulations from the joint model $\pi(\theta,y)$, amortized inference tackles this problem by directly learning the evidence (aka. evidential learning; Radev et al., 2023a, ) or by simultaneously learning an amortized posterior and likelihood approximator (Radev et al., 2023b, ). By design, prior-based model comparison is sensitive to the choice of prior (Schad et al.,, 2023; Oelrich et al.,, 2020), which often leads to issues in practical scenarios where prior specification is a hard challenge (Aguilar and Bürkner,, 2023; Dellaportas et al.,, 2012; Van Dongen,, 2006; Zitzmann et al.,, 2021; Lindley,, 1957). Differently, many methods of posterior-based model comparison rely on uncertainty-aware model simulations from the posterior predictive distribution (Equation 13; for a review see Vehtari and Ojanen,, 2012).
@@ 5.4 Sensitivity analysis
Sensitivity analysis (aka. robustness analysis; Berger et al.,, 1994) constitutes a critical step in statistical modeling that focuses on understanding how the output of a model changes in response to variations in its inputs. These inputs can include parameters of the model, the specification of prior distributions and observation models in a Bayesian setting, or the choices made while pre-processing the data. As such, each analysis hides an iceberg of uncertainty (Wagenmakers et al.,, 2022), and sensitivity analysis is a principled approach to render this uncertainty tangible. In the PAD taxonomy (see Section 2), sensitivity analyses can target each of the three axes of Bayesian models: Different statistical models, varying posterior approximators, and different variations of the training data.
@@
Akin to tracing all sources of uncertainty in Bayesian analysis, keeping track of all possible choices in sensitivity analysis resembles a Garden of Forking Paths (Borges,, 1941; McElreath,, 2020). Taming this combinatorial explosion is the key to enabling large-scale sensitivity-analyses with finite computational resources. To this end, Kallioinen et al., (2023) developed a computationally efficient approach that uses importance sampling to estimate the effect of power-scaling the prior and likelihood in a Bayesian analysis. In the context of SBI with neural networks, Elsemüller et al., (2024) proposed sensitivity-aware amortized Bayesian inference, which extends amortized Bayesian inference by a principled technique to encode context information about the probabilistic model. By additionally conditioning on this context information during the simulation-based training stage, the trained neural networks can subsequently perform near-instant sensitivity analyses without worrying about the associated cost of modifying model assumptions during inference.

### v2

@@ 4.2 Parameter Estimation
Posterior distributions: For a given prior $\pi(\theta)$, we are interested in updating the prior to the posterior $\pi(\theta\mid y)$, which incorporates all information about $\theta$ carried by the data $y$,
@@
Needless to say, both problems are computationally challenging for non-analytic models (e.g., no closed-form for the Bayesian posterior or unknown frequentist sampling distribution). The situation is further aggravated by P models defined solely through a (randomized) algorithm for generating data $y$ from parameter configurations $\theta$ instead of explicitly assuming a parametric data model $\pi(y\mid\theta)$ (Diggle and Gratton,, 1984). For such implicit models, the construction of confidence sets is challenging not only because we cannot evaluate the likelihood, but also because we need to test null hypotheses across the entire parameter space (Dalmasso et al.,, 2024). By the same token, posterior approximation is difficult because not only the marginal likelihood $\pi(y)$ involves an intractable integral, but also the implicit likelihood (Cranmer et al.,, 2020), rendering posterior approximation doubly intractable.
@@
Despite their theoretical elegance (Frazier et al.,, 2018), most ABC methods share the fundamental limitation of all non-amortized methods: computations must be repeated from scratch each time a model is fit to new data. Yet, there are many scenarios in which multiple model refits are necessary—for example, when modeling multiple data sets (von Krause et al.,, 2022), performing in silico model verification (Section 3), or aiming for real-time inference (Zeng et al.,, 2025). To address these challenges, a different way of utilizing simulations was needed—a way to pool or “compile” computations into global estimators that can produce near-instant results for arbitrary queries (Gershman and Goodman,, 2014; Paige and Wood,, 2016; Le et al.,, 2017). In hindsight, neural networks, as modular and flexibly composable universal function approximators, proved to be the ideal choice.
@@
Amortized inference asks how to flexibly reuse inferences or estimators in order to answer numerous queries without recomputation overhead. It has been proposed as a model of human probabilistic reasoning (Gershman and Goodman,, 2014), a method for fast inversion of graphical models (Stuhlmüller et al.,, 2013), and a means of learning variational posteriors with neural networks (Kingma and Welling,, 2013; Rezende et al.,, 2014). For statistical inference with “white-box” models, the latter idea extends to using simulations as training data for neural networks that approximate quantities of interest, such as point estimates, test statistics, or full posterior distributions (Bürkner et al.,, 2023; Zammit-Mangion et al.,, 2024, see also Figure 5). Neural networks are particularly favorable for amortized inference if the statistical model is sufficiently high dimensional (in data and/or parameter space), no good hand-crafted summary statistics are available, and model simulations can be obtained in reasonable time (to provide sufficient training budget).
@@
$\widehat{\mathcal{R}}(y_{\text{obs}}):=\{\theta\in\Theta\mid\lambda(y_{\text{obs}};\theta)\geq\widehat{C}_{\alpha}\}.$ (8)
@@
$\mathcal{L}(q):=\mathbb{E}_{\pi(\theta,y)}\Big{[}\mathcal{S}(q(\cdot\mid y),\theta)\Big{]}\approx\frac{1}{M}\sum_{m=1}^{M}\mathcal{S}(q(\cdot\mid y^{(m)}),\theta^{(m)}),$ (9)
which would guarantee $q(\theta\mid y)=\pi(\theta\mid y)$ under perfect convergence for large simulation budgets $M\rightarrow\infty$ (Gneiting et al.,, 2007) and a universal density approximator $q$ (e.g., coupling-based normalizing flows, Draxler et al.,, 2024). For instance, using the log-score for $S$, we retrieve the popular maximum likelihood objective,
@@
Evaluating Equation 10 empirically requires generative networks $q(\theta\mid y)$ with tractable densities, such as normalizing flows. More recently, score-based diffusion models have entered ABI (Sharrock et al.,, 2022; Geffner et al.,, 2023; Gloeckler et al.,, 2024), aimed at estimating the score of the (log) posterior density $\nabla_{\theta}\log\pi(\theta\mid y)$ by minimizing a tractable re-formulation of the weighted Fisher divergence:
@@
where $w(t)$ is a positive weight function, $s$ is a neural network, and $\pi_{t}(\theta^{(t)},y)$ is the model distribution defined in diffusion time $t$. The most popular tractable version of Equation 11 is conditional denoising (Song et al.,, 2021), but the framework includes various other “free-form” model families capable of learning from simulations, such as flow matching (Albergo et al.,, 2023; Lipman et al.,, 2022).
Notably, it was not initially evident that neural networks could serve as standalone global posterior approximators. Early approaches, for example, combined neural networks with ABC (Blum and François,, 2010) or SMC (Paige and Wood,, 2016). Papamakarios and Murray, (2016, p. 3) noted that generative networks could be trained over the entire prior predictive distribution $\pi(y)$, but dismissed the approach as “grossly inefficient”.
Figure 6: Estimation methods along the Pareto front balance a trade-off between inference speed and accuracy guarantees (Li et al.,, 2024).
However, subsequent work (Gonçalves et al.,, 2020; Radev et al.,, 2020) showed that amortization over the prior (and even across different priors, Elsemüller et al.,, 2024), combined with learnable data embeddings on the fly, is in fact a viable route to fully Bayesian inference. These developments were fueled by increasingly powerful neural density estimators (e.g., coupling-based normalizing flows, Dinh et al.,, 2017; Ardizzone et al.,, 2018) and greater representational capacity through algorithmic alignment (Bloem-Reddy and Teh,, 2020; Xu et al.,, 2020). Thus, it is now clear that amortization is not wasteful but rather essential for efficiently generating model-based insights across fields as varied as cognitive science (von Krause et al.,, 2022), evolutionary dynamics (Avecilla et al.,, 2022), and astrophysics (Dax et al.,, 2025), among others.
Importantly, the probabilistic calibration of all amortized methods can be efficiently checked via SBC (see Section 3.1 and Section 5.2), since the cost of independent posterior sampling over multiple model simulations is amortized. Additionally, the closeness of observed data to the typical set of simulations can be assessed via out-of-distribution (OOD) detection in (embedded) data space (Schmitt et al., 2023a, ; Huang et al.,, 2023). This is particularly important since amortized posteriors estimated from OOD data (e.g., resulting from model misspecification) tend to be biased (Schmitt et al., 2023a, ), which constitutes one of the major limitations of these methods.
Amortized methods highlight a fruitful merger between scientific simulation and deep learning. Despite their advantages in big data and likelihood-free settings, they lack the guarantees of gold-standard MCMC as part of the standard Bayesian workflow (Gelman et al.,, 2020) and their accuracy can break down when asked to extrapolate beyond simulated data (Frazier et al.,, 2024; Schmitt et al., 2023a, ). Recent work seeks to address these limitations through amortized workflows that combine ABI with MCMC (Li et al.,, 2024, see Figure 6) and trustworthy approaches that train on both simulated and real data (Elsemüller et al.,, 2025; Mishra et al.,, 2025; Swierc et al.,, 2024; Huang et al.,, 2023).
@@ 5 Model Checking
Model checking examines the compatibility of a statistical model with observed data after inference. This data-conditional perspective contrasts with model verification (Section 3), which typically relies on simulated data (e.g., from the prior predictive distribution). In contrast, model checking uses the observed data $y_{\text{obs}}$ to assess inference validity for the case of interest. Simulations are particularly important for scaling model checking algorithms to complex, high-dimensional settings.
@@ 5.1 Predictive Checking
If the data generating process is formalized as a simulation program, we can generate synthetic data replications based on the inferred parameter values. Predictive checking (Gelman and Hill,, 2006) combines parameter estimation (inverse problem) with simulation (forward problem), thereby casting the inference results back into the data space.
@@
In the Bayesian framework, we can propagate all uncertainty from the posterior through the data model. The result is the posterior predictive distribution, which expresses the distribution of new data $y^{\prime}$ while accounting for all uncertainty in the system,
@@
$\theta^{(s)}\sim\pi(\theta\mid y_{\text{obs}}),\quad y^{\prime(s)}_{1},\ldots,y^{\prime(s)}_{N}\sim\pi(y\mid\theta^{(s)},y_{\text{obs}})\quad\text{for}\;s=1,\ldots,S,$ (14)
where each simulated data set $y^{\prime(s)}=\{y^{\prime(s)}_{1},\ldots,y^{\prime(s)}_{N}\}$ contains $N$ observations. The draws from the posterior predictive distribution represent uncertainty-aware replications of the observed data (see Figure 7 for an illustrative example), where all associated uncertainty is propagated through the forward simulation process. Akin to frequentist predictive checking, any discrepancy statistic can then be computed from the observed data and the distribution of replicated data. For example, one might check whether the observed maximum value or a certain autocorrelation is extreme relative to the distribution of those metrics across the simulated replicates. If the observed data exhibit features that would be very unlikely under the posterior predictive distribution, it indicates model misfit.
@@
Figure 7: Statistical modeling of global temperature data with three parameters. Based on the observed global temperature $y_{\text{obs}}$ (7(a)), Bayesian inference yields draws from the posterior distribution of the parameters $\theta$ (7(b)). These posterior draws can be re-inserted into the simulation program (i.e., the global temperature model $f:\theta\mapsto y$) to obtain draws from the posterior predictive distribution (7(c)). These synthetic data replications contain uncertainty propagated from the posterior.
@@ 5.2 Posterior SBC
As detailed in Section 3.1.1, simulation-based calibration (SBC) checks if the inference is well-calibrated for data simulated from the prior predictive distribution. However, after observing real data $y_{\text{obs}}$, it is often more relevant to assess the inference conditional on that specific data rather than the entire prior predictive space. To address this, Posterior SBC (Säilynoja et al.,, 2025) uses the principles of SBC to validate the model’s implementation and inference algorithm conditioned on the observed data by means of simulations. This approach first draws $S$ samples $\theta^{\prime(s)}\sim\pi(\theta\mid y_{\text{obs}})$ from the (approximate) posterior conditional on the observed data $y_{\text{obs}}$. For each posterior draw, we then simulate data replications $y^{\prime(s)}$ from the posterior predictive distribution according to Equation 14. Finally, for each simulated data replication $y^{\prime(s)}$, we draw $D$ augmented posterior samples $\theta^{\prime\prime(s,1)},\ldots,\theta^{\prime\prime(s,D)}$ conditional on the observed data $y_{\text{obs}}$ and the respective data replication $y^{\prime(s)}$,
@@
and finally check whether the augmented posterior $\pi(\theta\mid y_{\text{obs}},y^{\prime})$ is well-calibrated with the “old” posterior $\pi(\theta\mid y_{\text{obs}})$ serving as prior. Pursuing a similar data-driven perspective, Fazio et al., (2024) propose adding small portions of real or simulated data to improve the robustness of Bayesian simulations, which in turn leads to increased stability for SBC and model simulations more broadly.
@@ 5.3 Model comparison
where $\pi(M)$ is a discrete prior distribution over the candidate models. In prior-based model comparison, the central quantity is the marginal likelihood $\pi(y_{\text{obs}}\mid M_{l})=\int\pi(y_{\text{obs}}\mid\theta,M_{l})\,\pi(\theta\mid M_{l})\mathrm{d}\theta$, which quantifies the evidence for model $M_{l}$ by integrating over all parameter values. This enables the computation of Bayes Factors (Kass and Raftery,, 1995) and posterior model probabilities. Since computing the marginal likelihood requires solving a potentially high-dimensional integral, this family of model comparison methods is typically computationally infeasible for complex models. Based on synthetic simulations from the joint model $\pi(\theta,y)$, amortized inference tackles this problem by directly learning the evidence (aka. evidential learning; Radev et al.,, 2021) or by simultaneously learning an amortized posterior and likelihood approximator (Radev et al.,, 2023).
By design, prior-based model comparison is sensitive to the choice of prior (Schad et al.,, 2023; Oelrich et al.,, 2020), which often leads to issues in practical scenarios where prior specification is a hard challenge (Aguilar and Bürkner,, 2023; Dellaportas et al.,, 2012; Van Dongen,, 2006; Zitzmann et al.,, 2021; Lindley,, 1957). Differently, posterior-based methods, such as cross-validation (CV), estimate the expected predictive distribution over new data, $\mathbb{E}\left[\pi(y^{\prime}\mid y_{\text{obs}},M_{l})\right]$ (Vehtari et al.,, 2017, 2019). Taken to the extreme, leave-one-out cross-validation (LOO-CV) evaluates the predictive distribution of single held-out observations (Vehtari et al.,, 2017). Importantly, amortized methods can perform overhead-free LOO-CV (or other cross-validation schemes) in cases where established approximate methods (e.g., Pareto-smoothed importance sampling, Vehtari et al.,, 2024) are infeasible (Radev et al.,, 2023).
@@ 5.4 Sensitivity analysis
Sensitivity analysis (aka. robustness analysis; Berger et al.,, 1994) constitutes a critical step in statistical modeling that focuses on understanding how the output of a model changes in response to variations in its inputs. These inputs can include fixed (hyper-)parameters of the model, the specification of prior distributions and observation models in a Bayesian setting, or the choices made while pre-processing the data. As such, each analysis hides an iceberg of uncertainty (Wagenmakers et al.,, 2022), and sensitivity analysis is a principled approach to render this uncertainty tangible. In the PAD taxonomy (see Section 2), sensitivity analyses can target each of the three axes of Bayesian models: Different statistical models, varying posterior approximators, and different variations of the training data.
@@
Akin to tracing all sources of uncertainty in Bayesian analysis, keeping track of all possible choices in sensitivity analysis resembles a Garden of Forking Paths (Borges,, 1941; McElreath,, 2020). Taming this combinatorial explosion is the key to enabling large-scale sensitivity-analyses with finite computational resources. To this end, Kallioinen et al., (2023) developed a computationally efficient approach that uses importance sampling to estimate the effect of power-scaling the prior and likelihood in a Bayesian analysis. Elsemüller et al., (2024) proposed sensitivity-aware amortized Bayesian inference, which extends the scope of amortization by encoding context information about the statistical model. By additionally conditioning on this context information during the simulation-based training stage, the trained neural networks can subsequently perform near-instant sensitivity analyses without worrying about the associated cost of modifying model assumptions during inference.
5.5 Increasing the efficiency of simulation-based estimators
Many use cases of simulations we highlighted above apply the standard Monte Carlo (MC) estimator $S^{-1}\sum_{s=1}f(\theta^{(s)})$ for approximating a target expectation $\mathbb{E}[f(\theta)]$. However, simulation-based estimators can become more sample-efficient than the standard MC estimator if additional information is available and correctly utilized. For example, in latent variable models, analytically integrating out the latent variables (if possible) before performing predictive simulations on new data can increase the stability of the resulting estimators (Merkle et al.,, 2019). As another example, “control variates” can improve the efficiency of Monte Carlo estimators by leveraging the score (i.e., the gradient of the log target density) when available (South et al.,, 2023). More generally, if conditional expectations of the form $\mathbb{E}[f(\theta)\mid T(y)]$ for an informative statistic $T(y)$ are available, approximating these conditional expectations via simulations can yield a more efficient estimator of $\mathbb{E}[f(\theta)]$, a process also known as Rao-Blackwellization (see Robert and Roberts,, 2021, for an overview). These examples illustrate that simulation-based methods are not an orthogonal alternative to analytical or density-based methods but can also benefit from the latter provided the right type of additional information.